\begin{thebibliography}{}
\bibitem{efetov-1}
%arXiv:cond-mat/0606620
%doi10.1103/PhysRevLett.98.016802
P. G. Silvestrov, K. B. Efetov,
%Quantum dots in graphene,
{\it Phys.Rev.Lett.} {\bf 98} (2007) 016802.
%Abstract: We suggest a way of confining quasiparticles by an external potential in a small region of a graphene strip. Transversal electron motion plays a crucial role in this %confinement. Properties of thus obtained graphene quantum dots are investigated theoretically for different types of the boundary conditions at the edges of the strip. The %(quasi)bound states exist in all systems considered. At the… ▽ More
\bibitem{efetov-2}
%arXiv:0707.3463
%doi10.1103/PhysRevB.77.155436
P. G. Silvestrov, K. B. Efetov,
%Charge accumulation at the boundaries of a graphene strip induced by a gate voltage: Electrostatic approach,
{\it Phys.Rev.} {\bf B77} (2008) 155436.
%Abstract: Distribution of charge induced by a gate voltage in a graphene strip is investigated. We calculate analytically the charge profile and demonstrate a %strong(macroscopic) charge accumulation along the boundaries of a micrometers-wide strip. This charge inhomogeneity is especially important in the quantum Hall regime where we %predict the doubling of the number of edge states and coexistence of two… ▽ More
\bibitem{matulis}
A. Matulis, F. M. Peeters,
%Quasibound states of quantum dots in single and bilayer graphene,
{\it Phys.Rev.} {\bf B77} (2008) 115423.
\bibitem{jakubsky-2}
%arXiv:1306.2115
%doi10.1088/1751-8113/47/11/115307
V. Jakubsky, S. Kuru, J. Negro,
%Carbon nanotubes in almost homogeneous transverse magnetic field: exactly solvable model,
{\it J.Phys.A} {\bf A47} (2014) 115307.
%Abstract: A class of exactly solvable models describing carbon nanotubes in the presence of an external inhomogeneous magnetic field is considered. The framework of the %continuum approximation is employed, where the motion of the charge carriers is governed by the Dirac- Weyl equation. The explicit solution of a particular example is provided. %It is shown that these models possess nontrivial integrals of mo…
\bibitem{jakubsky-3}
%arXiv:1405.2535 [pdf, ps, other] cond-mat.mes-hall
%doi10.1016/j.aop.2014.06.020
V. Jakubsky, D. Krejcirik,
%Qualitative analysis of trapped Dirac fermions in graphene,
{\it Ann.Phys.} {\bf 349} (2014) 268.
%Abstract: We study the confinement of Dirac fermions in graphene and in carbon nanotubes by an external magnetic field, mechanical deformations or inhomogeneities in the %substrate. By applying variational principles to the square of the Dirac operator, we obtain sufficient and necessary conditions for confinement of the quasi-particles. The %rigorous theoretical results are illustrated on the realistic examp… ▽ More
%Submitted 7 January, 2015; v1 submitted 11 May, 2014; originally announced May 2014.
\bibitem{jakubsky-1}
%arXiv:1412.1026
%doi10.1103/PhysRevD.91.045039
V. Jakubsky,
%Spectrally isomorphic Dirac systems: graphene in electromagnetic field,
{\it Phys.Rev.} {\bf D91} (2015) 045039.
%Abstract: We construct the new one-dimensional Dirac Hamiltonians that are spectrally isomorphic (not isospectral) with the known exactly solvable models. Explicit formulas for %their spectra and eigenstates are provided. The operators are utilized for description of Dirac fermions in graphene in presence of an inhomogeneous electromagnetic field. We %discuss explicit, physically relevant, examples of spectra… ▽ More
\bibitem{IN}
% Zero Energy States for a Class of Two-Dimensional Potentials in Graphene
M. V. Ioffe, D. N. Nishnianidze,
%Abstract: The excitations in graphene and some other materials are described by two-dimensional massless Dirac equation with applied external potential of %some kind. Solutions of this zero energy equation are built analytically for a wide class of scalar potentials. In contrast to most publications on %analytical solutions of massless two-dimensional Dirac equation, our potentials really depend on both spati… ▽ More
{\it Mod.Phys.Lett.} {\bf B32} (2018) 1850329. 
%\bibitem{fernandez-1}
%arXiv:1703.06190
%E. Díaz-Bautista, D. J. Fernandez C.,
%Graphene coherent states,
%{\it Eur.Phys.J.Plus} {\bf 132} (2017) 499.
%Abstract: In this paper, we will construct the coherent states for a Dirac electron in graphene placed in a constant homogeneous magnetic field which is orthogonal to the %graphene surface. First of all, we will identify the appropriate annihilation and creation operators. Then, we will derive the coherent states as eigenstates of the annihilation %operator, with complex eigenvalues. Several physical quantiti… ▽ More
\bibitem{novoselov-2}
K. S. Novoselov et. al.,
%A.K.Geim, S.V.Morozov, D.Jiang, M.I.Katsnelson, I.V.Grigorieva, S.V.Dubonos, and A.A.Firsov,
%Two-dimensional gas of massless Dirac fermions in graphene,
{\it Nature} {\bf 438} (2005) 197.
%Abstract
%Quantum electrodynamics (resulting from the merger of quantum mechanics and relativity theory) has provided a clear understanding of phenomena ranging from particle physics to %cosmology and from astrophysics to quantum chemistry. The ideas underlying quantum electrodynamics also influence the theory of condensed matter, but quantum relativistic %effects are usually minute in the known experimental systems that can be described accurately by the non-relativistic Schrödinger equation. Here we report an experimental study %of a condensed-matter system (graphene, a single atomic layer of carbon) in which electron transport is essentially governed by Dirac's (relativistic) equation. The charge %carriers in graphene mimic relativistic particles with zero rest mass and have an effective 'speed of light' c* approximately 10(6) m s(-1). Our study reveals a variety of %unusual phenomena that are characteristic of two-dimensional Dirac fermions. In particular we have observed the following: first, graphene's conductivity never falls below a %minimum value corresponding to the quantum unit of conductance, even when concentrations of charge carriers tend to zero; second, the integer quantum Hall effect in graphene is %anomalous in that it occurs at half-integer filling factors; and third, the cyclotron mass m(c) of massless carriers in graphene is described by E = m(c)c*2. This %two-dimensional system is not only interesting in itself but also allows access to the subtle and rich physics of quantum electrodynamics in a bench-top experiment.
\bibitem{novoselov}
A. N. Castro et. al.,
%The electronic properties of grapheme,
{\it Rev.Mod.Phys.} {\bf 81} (2009) 109.
%-162.
\bibitem{advances}
D. S. I. Abergel et. al.,
%Properties of graphene: a theoretical perspective,
{\it Advances in Physics} {\bf 59} (2010) 261.
%-482.
\bibitem{kats}
%arXiv:cond-mat/0612534
M. I. Katsnelson,
%Graphene: carbon in two dimensions,
%Abstract: Carbon is one of the most intriguing elements in the Periodic Table. It forms many allotropes, some being known from ancient times (diamond and graphite) and some %discovered ten to twenty years ago (fullerenes, nanotubes). Quite interestingly, the two-dimensional form (graphene) has been obtained only very recently, and immediately %attracted great deal of attention. Electrons in graphene, obeyin… ▽ More
{\it Materials Today} {\bf 10} (2007) 20.
%-27.
\bibitem{cooper}
F.Cooper, A.Khare, U.Sukhatme, {\it Phys.Rep.} {\bf 251} (1995) 267.
\bibitem{david}
D.J.Fernandez C., {\it AIP Conf.Proc.} {\bf 1287} (2010) 3.
\bibitem{AI}
A.A.Andrianov, M.V.Ioffe, {\it J.Phys.} {\bf A45} (2012) 503001.
\bibitem{peres}
%arXiv:0709.0646
%doi10.1088/0953-8984/19/40/406231
N. M. R. Peres, E. V. Castro,
%Algebraic solution of a graphene layer in a transverse electric and perpendicular magnetic fields,
{\it J.Phys.:Cond.Matt.} {\bf 19} (2007) 406231.
%Abstract: We present an exact algebraic solution of a single graphene plane in transverse electric and perpendicular magnetic fields. The method presented gives both the %eigen-values and the eigen-functions of the graphene plane. It is shown that the eigen-states of the problem can be casted in terms of coherent states, which appears in a %natural way from the formalism. ▽ More
\bibitem{kuru}
S. Kuru, J. Negro, L. M. Nieto,
%%Exact analytic solutions for a Dirac electron moving in graphene under magnetic fields,
{\it J.Phys.:Cond.Matt.} {\bf 21} (2009) 455305.
\bibitem{bardarson}
%arXiv:0902.4499
%doi10.1103/PhysRevLett.102.226803
J. H. Bardarson, M. Titov, P. W. Brouwer,
%Electrostatic confinement of electrons in an integrable graphene quantum dot,
{\it Phys.Rev.Lett.} {\bf 102} (2009) 226803.
%Abstract: We compare the conductance of an undoped graphene sheet with a small region subject to an electrostatic gate potential for the cases that the dynamics in the gated %region is regular (disc-shaped region) and classically chaotic (stadium). For the disc, we find sharp resonances that narrow upon reducing the area fraction of the gated region. %We relate this observation to the existence of confined… ▽ More
\bibitem{portnoi-4}
%arXiv:0908.0561
%doi10.1103/PhysRevB.81.245431
R. R. Hartmann, N. J. Robinson, M. E. Portnoi,
%Smooth electron waveguides in graphene,
{\it Phys.Rev.} {\bf B81} (2010) 245431.
%Abstract: We present exact analytical solutions for the zero-energy modes of two-dimensional massless Dirac fermions fully confined within a smooth one-dimensional potential %V(x)= - α/cosh(βx), which provides a good fit for potential profiles of existing top-gated graphene structures. We show that there is a threshold value of the characteristic %potential strength α/β for which the first mode appears, in st… ▽ More
%\bibitem{milpas}
%E. Milpas, M. Torres, G. Murguía,
%Magnetic field barriers in graphene: an analytically solvable model,
%{\it Condensed Matter} {\bf 23} (2011) 245304.
\bibitem{portnoi-1}
%arXiv:1105.0891
%doi10.1103/PhysRevB.84.155437
C. A. Downing, D. A. Stone, M. E. Portnoi,
%Zero-energy states in graphene quantum dots and rings,
{\it Phys.Rev.} {\bf B84} (2011) 155437.
%Abstract: We present exact analytical zero-energy solutions for a class of smooth decaying potentials, showing that the full confinement of charge carriers in electrostatic %potentials in graphene quantum dots and rings is indeed possible without recourse to magnetic fields. These exact solutions allow us to draw conclusions on the general %requirements for the potential to support fully confined states, incl… ▽ More
\bibitem{fernandez-2}
%arXiv:1402.4584
%doi10.1088/1751-8113/47/28/285302
B. Midya, D. J. Fernandez C.,
%Dirac electron in graphene under supersymmetry generated magnetic fields,
{\it J.Phys.} {\bf A47} (2014) 285302.
%Abstract: We use supersymmetry transformations to obtain new one parameter family of inhomogeneous magnetic fields B for which the massless Dirac electron possesses exact %solution. The inhomogeneity appearing in B˜(x,λ) can be controlled by the parameter λ. The obtained magnetic fields are interpreted as deformed variants of some p… ▽ More
\bibitem{portnoi-2}
%arXiv:1305.4652
%doi10.1103/PhysRevA.89.012101
R. R. Hartmann, M. E. Portnoi,
%Quasi-exact solution to the Dirac equation for the hyperbolic secant potential,
{\it Phys.Rev.} {\bf A89} (2014) 012101.
%Abstract: We analyze bound modes of two-dimensional massless Dirac fermions confined within a hyperbolic secant potential, which provides a good fit for potential profiles of %existing top-gated graphene structures. We show that bound states of both positive and negative energies exist in the energy spectrum and that there is a threshold value of the %characteristic potential strength for which the first mode… ▽ More
\bibitem{ghosh}
%arXiv:1512.02841
%doi10.1016/j.physleta.2015.12.006
P. Ghosh, P. Roy,
%An analysis of the zero energy states in graphene,
{\it Phys.Lett.} {\bf A380} (2015) 567.
%-569.
%Abstract: Using the concept of complex non PT symmetric potential we study creation of zero energy states in graphene by a scalar potential. The admissible range of the %potential parameter values for which such states exist has been examined. The situation with respect to the holes has also been investigated. ▽ More
\bibitem{ho-2}
%arXiv:1507.02649
%doi10.1209/0295-5075/112/47004
C.-L. Ho, P. Roy,
%mKdV equation approach to zero energy states of graphene,
{\it Eur.Phys.Lett.} {\bf 112} (2015) 47004.
%Abstract: We utilize the relation between soliton solutions of the mKdV and the combined mKdV-KdV equation and the Dirac equation to construct electrostatic fields which yield %exact zero energy states of graphene. ▽ More
\bibitem{portnoi-3}
%arXiv:1608.01522
%doi10.1103/PhysRevB.94.045430
C. A. Downing, M. E. Portnoi,
%Magnetic quantum dots and rings in two dimensions,
{\it Phys.Rev.} {\bf B94} (2016) 045430.
%Abstract: We consider the motion of electrons confined to a two dimensional plane with an externally applied perpendicular inhomogeneous magnetic field, both with and without a %Coulomb potential. We find that as long as the magnetic field is slowly-decaying, bound states in magnetic quantum dots are indeed possible. Several example cases of such %magnetic quantum dots are considered in which one can find the… ▽ More
\bibitem{portnoi-11}
C. A. Downing, M. E. Portnoi,
%Magnetic quantum dots and rings in two dimensions,
{\it Phys.Rev.} {\bf B94} (2016) 165407.
\bibitem{portnoi-111}
C. A. Downing, M. E. Portnoi,
%Magnetic quantum dots and rings in two dimensions,
{\it J.Phys.: Cond.Matt.} {\bf 29} (2017) 315301.
\bibitem{2-part}
C. A. Downing, M. E. Portnoi,
%Magnetic quantum dots and rings in two dimensions,
{\it Nature Communications} {\bf 8} (2017) 897.
\bibitem{midya}
B.Midya, D.J.Fernandez C., {\it J.Phys.} {\bf A47} (2014) 285302.
\bibitem{roy}
A.Schulze-Halberg, P.Roy, {\it J.Phys.} {\bf A50} (2017) 365205.
\bibitem{ho-1}
%arXiv:1405.6961
%doi10.1209/0295-5075/108/20004
C.-L. Ho, P. Roy,
%On zero energy states in graphene,
{\it Eur.Phys.Lett.} {\bf 108} (2014) 20004.
%Abstract: We obtain zero energy states in graphene for a number of potentials and discuss the relation of the decoupled Schrödinger-like equations for the the spinor components %with non relativistic PT symmetric quantum mechanics. ▽ More
\bibitem{setare}
D.Jahani, F.Shahbazy, M.R.Setare, {\it Eur.Phys.J.Plus} {\bf 133} (2018) 328.
\bibitem{schulze2019}
%Darboux transformations for the massless Dirac equation with matrix potential: Construction of zero-energy states
A.Schulze-Halberg, M.Ojel, {\it Eur.Phys.J.Plus} {\bf 134} (2019) 49.
\bibitem{schulze2021}
%First-order Darboux transformations for Dirac equations with arbitrary diagonal potential matrix in two dimensions
A.Schulze-Halberg, {\it Eur.Phys.J.Plus} {\bf 136} (2021) 790.
\bibitem{kuru2023}
%Graphene Dirac fermions in symmetric electric and magnetic fields: the case of an electric square well
I.B.Ateş, Ş.Kuru, J.Negro,
{\it Physica Scripta} {\bf 98} (2023) 015816.
\bibitem{INP}
M.V.Ioffe, D.N.Nishnianidze, E.V.Prokhvatilov,
{\it Eur.Phys.J.Plus} {\bf 134} (2019) 450.
\bibitem{jakubsky2020}
A.Contreras-Astorga, F.Correa, V.Jakubsky, {\it Phys.Rev.} {\bf B102} (2020) 115429.
\bibitem{IN2022}
M.V.Ioffe, D.N.Nishnianidze, {\it Eur.Phys.J.Plus} {\bf 137} (2022) 1195.
\bibitem{fokker}
M.V.Ioffe, D.N.Nishnianidze, {\it Eur.Phys.Lett.} {\bf 129} (2020) 61001.
%\bibitem{AIS}
%A.A.Andrianov, M.V.Ioffe, V.P.Spiridonov, {\it Phys.Lett.} {\bf A174} (1993) 273.
%\bibitem{ACDI}
%A.A.Andrianov, F.Cannata, J.P.Dedonder, M.V.Ioffe, {\it Int.J.Mod.Phys.} {\bf A10} (1995) 2683.
\bibitem{new}
F.Cannata, M.V.Ioffe, D.N.Nishnianidze, {\it J.Phys.} {\bf A35} (2002) 1389.
\bibitem{polyanin}
A.D.Polyanin, V.F.Zaitsev, {\it Handbook of Exact Solutions of Ordinary Differential Equations} (CRC Press, 1995).
\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{psfrag}
\usepackage{graphicx}
\usepackage{psfrag}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\ba}{\begin{eqnarray}}
\newcommand{\ea}{\end{eqnarray}}

\begin{document}
\begin{center}
{\bf
Solvable Two-dimensional Dirac Equation with Matrix Potential: Graphene in External Electromagnetic Field.
}\\
\vspace{1cm}
{\large M. V. Iof\/fe$^{1,}$\footnote{E-mail: ioffe000@gmail.com; corresponding author},
D. N. Nishnianidze}$^{2,}$\footnote{E-mail: cutaisi@yahoo.com}\\
\vspace{0.5cm}
$^1$ Saint Petersburg State University, 7/9 Universitetskaya nab., St.Petersburg, 199034 Russia.\\
$^2$ Akaki Tsereteli State University, 4600 Kutaisi, Georgia.\\

\end{center}
\vspace{0.5cm}
\hspace*{0.5in}
\vspace{1cm}
\hspace*{0.5in}
\begin{minipage}{5.0in}
{\small
It is known that the excitations in graphene-like materials in external electromagnetic field are described by solutions of massless two-dimensional Dirac equation which includes both Hermitian off-diagonal matrix and scalar potentials. Up to now, such two-component wave functions were calculated for different forms of external potentials but, as a rule, depending on one spatial variable only. Here, we shall find analytically the solutions for a wide class of combinations of matrix and scalar external potentials which physically correspond to applied mutually orthogonal magnetic and longitudinal electrostatic fields, both depending really on two spatial variables. The main tool for this progress was provided by supersymmetrical (SUSY) intertwining relations, namely, by their most general - asymmetrical - form proposed recently by the authors. Such SUSY-like method is applied in two steps similarly to the second order factorizable (reducible) SUSY transformations in ordinary Quantum Mechanics.
}
\end{minipage}

Keywords: two-dimensional Dirac equation, SUSY intertwining relations, matrix potential, graphene in electromagnetic field

{\it PACS:} 03.65.-w; 73.22.Pr

\section{Introduction.}

The extensive study of two-dimensional massless Dirac equation in the presence of external electromagnetic fields \cite{efetov-1},
\cite{efetov-2}, \cite{matulis}, \cite{jakubsky-2}, \cite{jakubsky-3}, \cite{jakubsky-1}, \cite{IN} 
 is due to its connection with the properties of electron carriers in graphene and in graphene-like materials \cite{novoselov-2}, \cite{novoselov}, \cite{advances}, \cite{kats}. The actual task is to find analytically the normalizable solutions of such Dirac equation where two components of "spinor" wave function $\Psi (\vec x)\equiv (\Psi_A(\vec x),\,\Psi_B(\vec x))$ correspond to two sublattices of graphene. The potentials in the equation have different origin: the off-diagonal matrix term is provided by the electromagnetic vector-potential $\vec A =(A_1 (x_1, x_2), A_2 (x_1, x_2), 0)$ in the "long derivatives" and leads to the magnetic field $\vec B = \vec\nabla\times\vec A$ along the z-axis, while the scalar potential $A_0 (x_1, x_2)$ describes the interaction with electrostatic or some another scalar field:
\be
[\sigma_1(-i\partial_1-A_1(\vec x)) + \sigma_2(-i\partial_2-A_2(\vec x)) + A_0(\vec x)]\Psi(\vec x)=0,   \label{dirac}
\ee
where two-dimensional $\vec x\equiv (x_1,\, x_2),$ the derivatives $\partial_i\equiv \frac{\partial}{\partial x_i},$ and the charge was taken $e=1.$


It is clear that the case of pure magnetic field in massless two-dimensional Dirac equation is explicitly solvable. Indeed, multiplying (\ref{dirac}) without the term $A_0$ by $\sigma_1,$ one obtains a pair of decoupled first order equations which can be easily solved. The presence in (\ref{dirac}) of term proportional to the unity matrix $\sigma_0$ prevents this decoupling (analogous problem would appear also in presence of the mass term proportional to $\sigma_3$). Different methods were used \cite{peres} - \cite{2-part} to study such two-dimensional Dirac equation, mainly with strong restrictions on the conditions of the problem. There were problems with only electrostatic or only magnetic fields, and also the problems with different specific one-dimensional ansatzes for external fields depending on a variable $x_1,$ or radial variable $r.$ 

The supersymmetrical method inherent in Schr\"odinger Quantum Mechanics \cite{cooper}, \cite{david}, \cite{AI} has become also one of the most effective tools in discussed problems \cite{midya}, \cite{roy}, \cite{ho-1}, \cite{setare}, \cite{schulze2019},\cite{schulze2021},     \cite{kuru2023}. As a rule, the main ingredient of SUSY Quantum Mechanics - so called SUSY intertwining relations - was explored in different forms. In the present paper, a class of external electromagnetic fields will be chosen with both magnetic (matrix) and electrostatic (scalar) terms depend effectively on both spatial coordinates. Such a progress is possible due 
to using a particular form of this approach - so called asymmetric intertwining relations \cite{INP}, \cite{jakubsky2020}, \cite{IN2022}. Up to now, this technique was explored to study the massless two-dimensional Dirac equation with scalar potential and also the equation of Fokker-Planck \cite{fokker}. More specifically, the procedure will include two steps (see Section 2). At first, asymmetric intertwining relations will provide SUSY diagonalization of potential in Eq.(\ref{dirac}) with both electromagnetic and electrostatic two-dimensional terms (Section 3). At the second stage, an additional asymmetric intertwining will connect the Dirac operator with diagonal potential with its partner whose potential is also diagonal but with constant elements (Section 4). 
The solutions of such Dirac equation can be found analytically, and solutions of initial problem are built by applying intertwining operators of both used steps (Section 5). Actually, the whole procedure realizes asymmetric form of factorizable SUSY intertwining of second order. Such SUSY intertwinings, but in their standard (symmetric) form, are known 
\cite{AI} in the context of SUSY Quantum Mechanics with Schr\"odinger operator.

\section{Asymmetric intertwining for two-dimensional Dirac equation in electromagnetic field.}

We start with the asymmetric intertwining relations:
\be 
D_1 N_1=N_2 D_2 \label{int}
\ee
for a pair of two-dimensional massless Dirac operators of the form (\ref{dirac}) rewritten as general operators with Hermitian matrix potentials
\be
D_{1,2} \equiv -i\sigma_k\partial_k + V_{1,2}(\vec x); \quad k=1,2; \quad V_{i}(\vec x) = \left(
                                                                                    \begin{array}{cc}
                                                                                      v^{(i)}_{11}(\vec x) & v^{(i)}_{12}(\vec x) \\
                                                                                      v^{(i)}_{21}(\vec x) & v^{(i)}_{22}(\vec x) \\
                                                                                    \end{array}
                                                                                  \right); \quad i=1,2. 
                                                                                  \label{D}
\ee
Here, two different intertwining operators have the general matrix form:
\be 
N_1 = A_k\partial_k + A(\vec x); \quad N_2 = B_k\partial_k + B(\vec x) \label{NM}
\ee
with constant matrices $A_k, \, B_k$ and two $\vec x-$dependent matrices $A(\vec x), B(\vec x):$
\be 
A(\vec x)=\left(
            \begin{array}{cc}
              a_{11}(\vec x) & a_{12}(\vec x) \\
              a_{21}(\vec x) & a_{22}(\vec x) \\
            \end{array}
          \right); \quad B(\vec x)=\left(
            \begin{array}{cc}
              b_{11}(\vec x) & b_{12}(\vec x) \\
              b_{21}(\vec x) & b_{22}(\vec x) \\
            \end{array}
          \right).
          \label{AABB}
\ee

Expanding Eq.(\ref{int}) in powers of derivatives we obtain:
\ba 
&&\sigma_kA_n-B_k\sigma_n+\sigma_nA_k-B_n\sigma_k=0,\quad k,n=1,2, \label{1-2}\\
&&i\sigma_kA(\vec x)-V_1(\vec x)A_k=-B_kV_2(\vec x)+iB(\vec x)\sigma_k,\quad k=1,2, \label{1-3}\\
&&i\sigma_k(\partial_kA(\vec x))-V_1(\vec x)A(\vec x)=-B_k(\partial_kV_2(\vec x))-B(\vec x)V_2(\vec x),\quad k=1,2. \label{1-4}
\ea
Equations (\ref{1-2}) provide the following form of constant coefficient matrices $A_k,\,B_k$:
\be 
A_1=\left(
      \begin{array}{cc}
        a & b \\
        d & c \\
      \end{array}
    \right); \quad A_2=\left(
      \begin{array}{cc}
        n & -ib \\
        id & p \\
      \end{array}
    \right); \quad B_1=\left(
      \begin{array}{cc}
        c & d \\
        b & a \\
      \end{array}
    \right); \quad B_2=\left(
      \begin{array}{cc}
        p & -id \\
        ib & n \\
      \end{array}
    \right); \label{AB}
\ee
with constant values $a,b,c,d,n,p,$ and Eqs.(\ref{1-3}) give the system of eight linear equations for the matrix elements of $V_{1,2}(\vec x), A(\vec x), B(\vec x):$
\ba  
&&(p+ic)(v^{(1)}_{12}-v^{(2)}_{12})=0; \label{3-2}\\
&&(n-ia)(v^{(1)}_{21}-v^{(2)}_{21})=0; \label{4-3}\\
&&(n+ia)v^{(1)}_{11}+2idv^{(1)}_{12}-(p+ic)v^{(2)}_{11}=2b_{12}; \label{3-1}\\
&&(n+ia)(v^{(1)}_{21}-v^{(2)}_{21})+2idv^{(1)}_{22}-2ibv^{(2)}_{11}=-2a_{11}+2b_{22}; \label{3-3}\\
&&(p+ic)v^{(1)}_{22}-2ibv^{(2)}_{12}-(n+ia)v^{(2)}_{22}=-2a_{12}; \label{3-4}\\
&&(n-ia)v^{(1)}_{11}+2idv^{(2)}_{21}-(p-ic)v^{(2)}_{11}=2a_{21}; \label{4-1}\\
&&(p-ic)(v^{(1)}_{12}-v^{(2)}_{12})+2idv^{(2)}_{22}-2ibv^{(1)}_{11}=-2b_{11}+2a_{22}; \label{4-2}\\
&&(p-ic)v^{(1)}_{22}-(n-ia)v^{(2)}_{22}-2ibv^{(1)}_{21}=-2b_{21}.  \label{4-4}
\ea 
The first two equations of the system indicates the need to highlight four different possibilities for solution of (\ref{3-2}) -- (\ref{4-4}) with Hermiticity of both potentials in mind:
\ba 
I) &&p+ic=n-ia=0; \nonumber\\
II) &&v^{(1)}_{12}=v^{(2)}_{12}; \quad v^{(1)}_{21}=v^{(2)}_{21}; \quad (p+ic)(n-ia) \neq 0\nonumber\\
III) &&v^{(1)}_{12}=v^{(2)}_{12}; \quad v^{(1)}_{21}=v^{(2)}_{21};\quad p+ic=0; \nonumber\\
IV) &&v^{(1)}_{12}=v^{(2)}_{12}; \quad v^{(1)}_{21}=v^{(2)}_{21};\quad n-ia=0. \nonumber
\ea 

Thus, Eqs.(\ref{1-2}), (\ref{1-3}) are reduced to the system of Eqs.(\ref{3-1}) - (\ref{4-4}) in four possible variants above. We have not considered yet the matrix differential equation (\ref{1-4}). It is convenient to solve it in an indirect form by using its combination with derivative $\partial_k$ of equation (\ref{1-3}):
\be 
V_1(\vec x)A(\vec x) - B(\vec x)V_2(\vec x)=(\partial_kV_1(\vec x))A_k+i(\partial_kB(\vec x))\sigma_k.
\label{1-9}
\ee

Below, all four variants will be used to obtain the general solution of initial intertwining relations (\ref{int}).

\section{SUSY-diagonalization by means of intertwining.}

Substitution of $p+ic=n-ia=0,$ i.e. of Variant I above, into other equations (\ref{3-1}) - (\ref{4-4}) together with Eqs.(\ref{1-9}) is quite long, but straightforward. These calculations lead to a system of four nonlinear equations, two of them being differential:
\ba 
v^{(1)}_{11}(ia_{11}+bv^{(2)}_{11}+av^{(2)}_{21})&=&v^{(2)}_{11}(ia_{22}+cv^{(2)}_{12}+dv^{(2)}_{22}); \label{10-1-1}\\
v^{(1)}_{22}(ia_{22}+cv^{(2)}_{12}+dv^{(2)}_{22})&=&v^{(2)}_{22}(ia_{11}+bv^{(2)}_{11}+av^{(2)}_{21}); \label{10-1-2}\\
i(v^{(2)}_{12}-v^{(1)}_{12}) (ia_{22}+cv^{(2)}_{12}+dv^{(2)}_{22})&=&2\partial_z(ia_{22}+cv^{(2)}_{12}+dv^{(2)}_{22}); \label{10-2-1}\\
i(v^{(2)}_{21}-v^{(1)}_{21}) (ia_{11}+bv^{(2)}_{11}+av^{(2)}_{21})&=&2\partial_{\bar z}(ia_{11}+bv^{(2)}_{11}+av^{(2)}_{21}). \label{10-2-2}
\ea 
Here and below, it is convenient to use the space arguments $\vec x$ of the functions in the form of complex variables $z=x_1+ix_2; \quad \bar z=x_1-ix_2$ and corresponding derivatives $\partial\equiv \frac{1}{2}(\partial_1-i\partial_2);\,\,\bar\partial\equiv \frac{1}{2}(\partial_1+i\partial_2)$. In particular, the system (\ref{10-1-1}) - (\ref{10-2-2}) takes the compact form:
\ba 
 v^{(1)}_{11}(z, \bar z) f_1(z, \bar z) &=& v^{(2)}_{11}(z, \bar z) f_2(z, \bar z); \label{11-1}\\
v^{(1)}_{22}(z, \bar z) f_2(z, \bar z) &=& v^{(2)}_{22}(z, \bar z) f_1(z, \bar z); \label{11-2}\\
v^{(1)}_{12}(z, \bar z)&=&v^{(2)}_{12}(z, \bar z)+2i\partial\ln (f_2(z,\bar z)); \label{11-3}\\
v^{(1)}_{21}(z, \bar z)&=&v^{(2)}_{21}(z, \bar z)+2i\bar\partial\ln (f_1(z,\bar z)), \label{11-4}
\ea 
where two combinations are introduced:
\ba 
f_1(z, \bar z)\equiv ia_{11}(z, \bar z)+bv^{(2)}_{11}(z,\bar z)+av^{(2)}_{21}(z, \bar z); \label{11-0}\\
f_2(z, \bar z)\equiv ia_{22}(z, \bar z)+cv^{(2)}_{12}(z,\bar z)+dv^{(2)}_{22}(z, \bar z). \label{11-0-0}
\ea 

One may notice the apparent paradox contained in the system (\ref{10-1-1}) - (\ref{10-2-2}). All these equations are identically fulfilled if
$$ia_{11}+bv^{(2)}_{11}+av^{(2)}_{21}=ia_{22}+cv^{(2)}_{12}+dv^{(2)}_{22}=0$$
for arbitrary potentials $V_1,\, V_2.$ The explanation is rather simple: just in this case, Dirac operators $D_1, D_2$ are proportional to the intertwining operators $D_2=CN_1, D_1=N_2C$ up to some constant matrix $C.$ Therefore, intertwining relation (\ref{int}) becomes a trivial identity in such a case.

In the present context, the typical way \cite{cooper}, \cite{david}, \cite{AI}, \cite{INP}, \cite{IN2022} of using SUSY intertwining relations can be formulated as follows. Let us choose one of the Dirac operators, $D_2,$ such that the corresponding Dirac equation is rather simple so that the problem of its analytical solution is more easy. Then, the solutions of partner Dirac equation with operator $D_1$ will be found by the action of intertwining operator $\Psi^{(1)}=N_1\Psi^{(2)}$ (in its turn, the operator $N_2^{\dagger}$ transforms the spinor $\Psi^{(1)}$ into $\Psi^{(2)}$). As the first step on this way, we choose the potential $V_2(\vec x)$ as a diagonal matrix:
\be 
V_2(\vec x) = diag(v^{(2)}_{11}(z,\bar z),\,\, v^{(2)}_{22}(z,\bar z))\equiv diag(v_1(z,\bar z),\,\, v_2(z,\bar z))
\label{V_2}
\ee
with real diagonal matrix elements. Then, due to Eqs.(\ref{11-1}), (\ref{11-2}), fraction $f_2(z, \bar z)/f_1(z, \bar z)$ is a real function. Futhermore,  Eqs.(\ref{11-3}), (\ref{11-4}) and the Hermiticity of $V_1(\vec x),$ i.e. $v^{(1)}_{12}(z,\bar z)=v^{(1)\star}_{21}(z,\bar z),$ lead to the following restriction:
\be 
f_1(z, \bar z)f^{\star}_2(z, \bar z)=c
\label{12-2}
\ee
with an arbitrary real constant $c.$ Summarizing these results, potential $V_1$ is expressed in terms of components (\ref{V_2}), function $f_2(z, \bar z)$ and real constant $c:$
\be 
V_1(\vec x)=\left(
              \begin{array}{cc}
                \frac{|f_2|^2}{c}v_1 &  2i\partial\ln (f_2)\\ 
                -2i\bar\partial\ln (f_2^{\star}) & \frac{c}{|f_2|^2}v_2 \\
              \end{array}
            \right)= \left(
              \begin{array}{cc}
                \pm f^2v_1 &  2i(\partial\ln (f) +i\partial\varphi)\\ 
                -2i(\bar\partial\ln (f) -i\bar\partial\varphi )& \pm f^{-2}v_2 \\
              \end{array}
            \right),
\label{12-4}
\ee
where the function $f_2$ was parameterizing as:
\be 
f_2(z, \bar z)\equiv \sqrt{\pm c}f(z, \bar z)e^{i\varphi(z, \bar z)}. \label{12-12}
\ee
Above, the sign $\pm$ corresponds to cases $c > 0,\, c < 0,$ respectively, and $f(z, \bar z)$ is a positive function. This function 
connects elements of initial diagonal potential $V_2$ as:
\be 
v_2(z, \bar z) = f^4(z, \bar z)v_1(z, \bar z).
\label{13-2}
\ee

For the physical system with matrix potential $V_1$ describing "spin $1/2$ particle" in external electromagnetic field according to Eq.(\ref{dirac}), the diagonal elements of $V_1$ define the electrostatic potential
\be 
A_0(\vec x)=v^{(1)}_{11}=v^{(1)}_{22}=\pm (v_1 v_2)^{1/2}=\pm f^2(\vec x)v_1(\vec x),
\label{A_0}
\ee
while the off-diagonal terms define the magnetic field which is orthogonal to the plane $x_1,x_2$:
\be 
B_3(\vec x)=\vec\nabla\times\vec A(\vec x)=\bigtriangleup \ln(f(\vec x)).
\label{13-6}
\ee
The solutions of Dirac equation with such external fields could be obtained from the two-component solutions $\Psi^{(2)}(\vec x)$ of Dirac equation with diagonal potential $V_2$ by action of the intertwining operator $N.$ But the diagonal form of potential $V_2$ does not yet provide solvability of the corresponding Dirac problem analytically. Up to now, supersymmetric intertwining relations (\ref{int}) connected an initial Dirac problem with operator $D_1$ to Dirac problem with the potential $V_2,$ which is diagonal, and therefore, has chances to be solvable. This part of procedure can be called SUSY-diagonalization of two-dimensional Dirac problem with matrix potential (by analogy with SUSY separation of variables \cite{new}).

 \section{From diagonal potential to the constant one by means of intertwining.}
 
 Let us consider now the variant II with $(p+ic)(n-ia)\neq 0$ for solution of the system (\ref{3-2})  - (\ref{4-4}) in the context of the second step of our procedure. Namely, let us consider intertwining relations between two Dirac operators, both with diagonal potential:
 \be 
U_1(\vec x)=
  \left(
              \begin{array}{cc}
                v_{1}(\vec x) & 0\\ 
                0 & v_{2}(\vec x) \\
              \end{array}
            \right); 
            \quad U_2(\vec x)= \left(
              \begin{array}{cc}
                m_1 & 0\\ 
                0 & m_2 \\
              \end{array}
            \right),
\label{01-0}
\ee
with constant elements $m_1, m_2.$ Here, potential $U_1(\vec x)$ is identified with potential $V_2(\vec x)$ of the previous step, and the constant partner potential $U_2$ will provide solvability of the problem. The intertwining operators $N_1, \, N_2$ have the same general form (\ref{NM}), and the explicit expressions for matrices $A_k,\, B_k,\, A(\vec x),\, B(\vec x)$ will be found below by analytical solution of the system of equations which were obtained in Section 2. 

We shall consider the system of equations (\ref{3-2}) - (\ref{4-4}) sequentially. Eqs.(\ref{3-2}) and (\ref{4-3}) are fulfilled automatically. Eqs.(\ref{3-1}), (\ref{4-1}) and (\ref{3-4}), (\ref{4-4}) allow to express off-diagonal elements of $A(\vec x),\, B(\vec x)$ at (\ref{AABB}) in terms of $v_1(\vec x), \, v_2(\vec x):$
\ba 
a_{12}(\vec x)&=&-\frac{1}{2}(p+ic)v_2(\vec x) + \frac{1}{2}m_2(n+ia); \label{03-2}\\
a_{21}(\vec x)&=&\frac{1}{2}(n-ia)v_2(\vec x) - \frac{1}{2}m_1(p-ic); \label{03-1}\\
b_{12}(\vec x)&=&\frac{1}{2}(n+ia)v_1(\vec x) - \frac{1}{2}m_1(p+ic); \label{03-3}\\
a_{12}(\vec x)&=&-\frac{1}{2}(p-ic)v_2(\vec x) + \frac{1}{2}m_2(n-ia). \label{03-4}
\ea
Eqs.(\ref{4-2}) and (\ref{3-3}) for diagonal elements of $A(\vec x),\, B(\vec x)$ can be written as:
\ba 
b_{11}(\vec x)=a_{22}(\vec x)+ibv_1(\vec x)-idm_2; \label{04-5}\\
b_{22}(\vec x)=a_{11}(\vec x)+idv_2(\vec x)-ibm_1. \label{04-6}
\ea

Eq.(\ref{1-4}) in its initial form is convenient to write now in complex coordinates $z\, \bar z:$
\ba  
&&2i\Biggl[\left(
              \begin{array}{cc}
                0 & \partial\\ 
                \bar{\partial} & 0 \\
              \end{array}
            \right) - \left(
              \begin{array}{cc}
                v_{1}(z,\bar z) & 0\\ 
                0 & v_{2}(z,\bar z) \\
              \end{array}
            \right)\Biggr] \left(
              \begin{array}{cc}
                a_{11}(z,\bar z) & a_{12}(z,\bar z)\\ 
                a_{21}(z,\bar z) & a_{22}(z,\bar z) \\
              \end{array}
            \right) = \nonumber \\ 
            &&=-\left(
              \begin{array}{cc}
                b_{11}(z,\bar z) & b_{12}(z,\bar z)\\ 
                b_{21}(z,\bar z) & b_{22}(z,\bar z) \\
              \end{array}
            \right) \left(
              \begin{array}{cc}
                m_1 & 0\\ 
                0 & m_2 \\
              \end{array}
            \right),
\label{04-0}
\ea 
where $\partial\equiv\partial_z=\frac{1}{2}(\partial_1-i\partial_2),\,\,\bar{\partial}\equiv\partial_{\bar z}=\frac{1}{2}(\partial_1+i\partial_2).$
In components, matrix equation (\ref{04-0}) is equivalent to the system of linear first order differential equations:
\ba 
2i(\partial a_{21}(z, \bar z))-v_1(z, \bar z)a_{11}(z, \bar z)&=&-m_1b_{11}(z, \bar z); 
\label{04-1}\\
2i(\bar{\partial} a_{12}(z, \bar z))-v_2(z, \bar z)a_{22}(z, \bar z)&=&-m_2 b_{22}(z, \bar z); 
\label{04-2}\\
2i(\partial a_{22}(z, \bar z))-v_1(z, \bar z)a_{12}(z, \bar z)&=&-m_2b_{12}(z, \bar z); 
\label{04-3}\\
2i(\bar{\partial} a_{11}(z, \bar z))-v_2(z, \bar z)a_{21}(z, \bar z)&=&-m_1b_{21}(z, \bar z). 
\label{04-4}
\ea
Let us define for convenience:
\be 
a_{11}(z, \bar z)-im_1b\equiv g_1(z, \bar z); \quad a_{22}(z, \bar z)-im_2d\equiv g_2(z, \bar z).
\label{06-0}
\ee
Then, after substitution of Eqs.(\ref{03-2}) - (\ref{03-4}) and Eqs.(\ref{04-5}), (\ref{04-6}), the system (\ref{04-1}) - (\ref{04-4}) takes the form:
\ba 
&&i(n-ia)(\partial v_{1}(z, \bar z))=v_1(z, \bar z)g_{1}(z, \bar z)-m_1g_{2}(z, \bar z); \label{06-1}\\
&&i(p+ic)(\bar{\partial} v_{2}(z, \bar z))=-v_2(z, \bar z)g_{2}(z, \bar z)+m_2g_{1}(z, \bar z); \label{06-2}\\
&&4i(\bar\partial g_{1}(z, \bar z))=(n-ia)(v_1(z, \bar z)v_2(z, \bar z)-m_1m_2); \label{06-3}\\
&&4i(\partial g_{2}(z, \bar z))=-(p+ic)(v_1(z, \bar z)v_2(z, \bar z)-m_1m_2), \label{06-4}
\ea
where the form of last two equations means that functions $g_1(z, \bar z),\, g_2(z, \bar z)$ can be expressed in terms of one complex function $g$:
\be 
g_1(z, \bar z)=-\frac{1}{p+ic}(\partial g(z, \bar z));\quad g_2(z, \bar z)=\frac{1}{n-ia}(\bar{\partial} g(z, \bar z)),
\label{06-34}
\ee
which satisfies the second order equation:
\be 
4i(\partial\bar{\partial}g(z, \bar z))=-(n-ia)(p+ic)(v_1(z, \bar z) v_2(z, \bar z)-m_1m_2).
\label{06-5}
\ee
and the first two equations (\ref{06-1}), (\ref{06-2}) become now:
\ba  
&&i(n-ia)(\partial v_{1}(z, \bar z))=-\frac{1}{p+ic}v_1(z, \bar z) (\partial g(z, \bar z))-\frac{1}{n-ia}m_1(\bar{\partial}g(z, \bar z)); \label{06-6}\\
&&i(p+ic)(\bar{\partial} v_{2}(z, \bar z))=-\frac{1}{n-ia}v_2(z, \bar z) (\bar{\partial} g(z, \bar z))-\frac{1}{p+ic}m_2(\partial g(z, \bar z)). \label{06-7}
\ea
Thus,
\ba 
&&\Omega (\partial v_1(z, \bar z)) + v_1(z, \bar z)(\partial g(z, \bar z)) = -L_1 (\bar{\partial} g(z, \bar z));
\label{07-2}\\
&&\Omega (\bar{\partial} v_2(z, \bar z)) + v_2(z, \bar z)(\bar{\partial} g(z, \bar z)) = -L_2 (\partial g(z, \bar z));
\label{07-3},\\
&&\Omega\equiv i(n-ia)(p+ic);\quad L_1\equiv\frac{m_1(p+ic)}{n-ia};\quad L_2\equiv\frac{m_2(n-ia)}{p+ic}\nonumber
\ea
Below, we shall solve this system of equation by considering separately different options for the choice of constant parameters.

\subsection{{\bf The case A:}  Parameters are real.}
 
Let us study the case with real function $g(z, \bar z)$ and real values of parameters $\Omega,\, L_1, \, L_2.$ Taking into account the reality 
of $v_1(z, \bar z)$ and $v_2(z, \bar z),$ it is useful to come back to Cartesian coordinates $x_1,\,x_2$ in Eqs.(\ref{07-2}) and (\ref{07-3}). Separately, both real and imaginary parts of (\ref{07-2}) are integrated explicitly with two "constants of integration" $s_1(x_1),\, s_2(x_2),$ which are arbitrary real functions of their arguments:
\ba 
\exp{(g(\vec x)/\Omega)} &=& L_1^{-1}(s_2(x_2)-s_1(x_1)); \label{08-11}\\
v_1(\vec x)&=&L_1\frac{s_2(x_2)+s_1(x_1)}{s_2(x_2)-s_1(x_1)}. \label{08-12}
\ea
Analogously, Eq.(\ref{07-3}) can be integrated as well with similar result:
\ba 
\exp{(g(\vec x)/\Omega)} &=& L_2^{-1}(\tilde{s}_2(x_2)-\tilde{s}_1(x_1)); \label{08-21}\\
v_2(\vec x)&=&L_2\frac{\tilde{s}_2(x_2)+\tilde{s}_1(x_1)}{\tilde{s}_2(x_2)-\tilde{s}_1(x_1)}, \label{08-22}
\ea
and arbitrary real $\tilde{s}_1(x_1), \,\tilde{s}_2(x_2).$
Eqs.(\ref{08-11}), (\ref{08-21}) together allow to connect $\tilde{s}_1(x_1), \,\tilde{s}_2(x_2)$ with their analogues:
\be 
\tilde{s}_1(x_1)=L_1^{-1}L_2(s_1(x_1)+\delta);\quad \tilde{s}_2(x_2)=L_1^{-1}L_2(s_2(x_2)+\delta),
\label{08-3}
\ee
with an arbitrary real constant $\delta .$ These relations have to be substituted into expression (\ref{08-22}).

Using these connections in second order differential equation (\ref{06-5}) for function $g(\vec x)$ and differentiating it by $\partial_1\partial_2,$ we obtain the simple third order equation with separable variables:
\be 
\frac{s_1^{\prime\prime\prime}(x_1)}{s_1^{\prime}(x_1)} + \frac{s_2^{\prime\prime\prime}(x_2)}{s_2^{\prime}(x_2)} = -4L_1L_2=-4m_1m_2.
\label{09-3}
\ee
After separation of variables in (\ref{09-3}) and integration of one-dimensional equations, we have:
\be 
s_1^{\prime\prime}(x_1)=\lambda_1^2s_1(x_1)+\omega_1\lambda_1^2; \quad   
s_2^{\prime\prime}(x_2)=\lambda_2^2s_2(x_2)+\omega_2\lambda_2^2,
\label{09-4}
\ee
where $\omega_1,\,\omega_2$ are integration constants, and $\lambda_1,\,\lambda_2$ - arbitrary constants which satisfy relation:
\be 
\lambda_1^2 + \lambda_2^2=-4m_1m_2.\nonumber 
\ee
Solutions of (\ref{09-4}) are known:
\ba 
s_1(x_1)&=&\frac{1}{2}(\sigma_1e^{\lambda_1x_1}+\delta_1e^{-\lambda_1x_1}) - \omega_1; \nonumber \\
s_2(x_2)&=&\frac{1}{2}(\sigma_2e^{\lambda_2x_2}+\delta_2e^{-\lambda_2x_2}) - \omega_2.\nonumber 
\ea
Because we used derivatives of Eq.(\ref{06-5}), it is necessary to check results. Substitution of (\ref{09-4}) into (\ref{06-5}) gives two relations between parameters:
\ba 
&&\omega_1\lambda_1^2 + \omega_2\lambda_2^2 + 2 \delta L_1L_2 = 0; \nonumber \\
&&\lambda_1^2(\sigma_1\delta_1-\omega_1^2) +\lambda_2^2(\sigma_2\delta_2-\omega_2^2)=0. \nonumber 
\ea

Now we can list, depending on the values of the constants, all possible solutions $s_1(x_1),\, s_2(x_2)$ within this Subsection. All of them are expressed in terms of hyperbolic, trigonometric and exponential functions, and they have to be inserted into (\ref{08-3}) to find $u_1(\vec x),\, u_2(\vec x)$ according to (\ref{08-12}), (\ref{08-22}):

{\bf I.}  $\lambda_1^2>0;\quad \lambda_2^2>0; \quad \sigma_1\delta_1>0; \quad \sigma_2\delta_2>0$
  
  By additional translation of $x_{1,2}$ functions $s_1(x_1),\, s_2(x_2)$ takes the form:
  \be 
  s_1(x_1)=\sigma_1\cosh(\lambda_1x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cosh(\lambda_2x_2)-\omega_2, \nonumber 
  \ee
  with restriction:
  \be 
  \lambda_1^2(\sigma_1^2-\omega_1^2) + \lambda_2^2(\sigma_2^2-\omega_2^2)=0. \nonumber 
  \ee
  
{\bf II.}  $\lambda_1^2>0;\quad \lambda_2^2>0; \quad \sigma_1\delta_1>0; \quad \sigma_2\delta_2<0$
  \be 
  s_1(x_1)=\sigma_1\cosh(\lambda_1x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\sinh(\lambda_2x_2)-\omega_2, \nonumber 
  \ee
  with restriction:
  \be 
  \lambda_1^2(\sigma_1^2-\omega_1^2) - \lambda_2^2(\sigma_2^2+\omega_2^2)=0.\nonumber 
  \ee
{\bf III.}  $\lambda_1^2>0;\quad \lambda_2^2>0; \quad \sigma_1\delta_1>0; \quad \delta_2=0$
  \be 
  s_1(x_1)=\sigma_1\cosh(\lambda_1x_1)-\omega_1;\quad s_2(x_2)=\frac{1}{2}\sigma_2\exp(\lambda_2x_2)-\omega_2, \nonumber 
  \ee
  with restriction:
  \be 
  \lambda_1^2(\sigma_1^2-\omega_1^2) - \lambda_2^2\omega_2^2=0. \nonumber 
  \ee

{\bf IV.}  $\lambda_1^2<0;\quad \lambda_2^2<0; \quad \sigma_1=\delta_1^{\star}; \quad \sigma_2=\delta_2^{\star}$
 
 $\lambda_1\equiv i\Lambda_1;\quad \lambda_2\equiv i\Lambda_2$
  \be 
  s_1(x_1)=\sigma_1\cos(\Lambda_1x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\Lambda_2x_2)-\omega_2; \nonumber 
  \ee
  with restriction:
  \be 
  \Lambda_1^2(\sigma_1^2-\omega_1^2) + \Lambda_2^2(\sigma_2^2-\omega_2^2)=0.\nonumber 
  \ee

{\bf V.}  $\lambda_1^2>0;\quad \lambda_2^2<0; \quad \sigma_1\delta_1>0; \quad \sigma_2=\delta_2^{\star}$
 
 $\lambda_2\equiv i\Lambda_2$
  \be 
  s_1(x_1)=\sigma_1\cosh(\lambda_1x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\Lambda_2x_2)-\omega_2; \nonumber 
  \ee
  with restriction:
  \be 
  \lambda_1^2(\sigma_1^2-\omega_1^2) - \Lambda_2^2(\sigma_2^2-\omega_2^2)=0. \nonumber 
  \ee
  
  {\bf VI.}  $\lambda_1^2>0;\quad \lambda_2^2<0; \quad \sigma_1\delta_1<0; \quad \sigma_2=\delta_2^{\star}$
 $\lambda_2\equiv i\Lambda_2$
  \be 
  s_1(x_1)=\sigma_1\sinh(\lambda_1x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\Lambda_2x_2)-\omega_2; \nonumber 
  \ee
  with restriction:
  \be 
  \lambda_1^2(\sigma_1^2+\omega_1^2) + \Lambda_2^2(\sigma_2^2-\omega_2^2)=0. \nonumber 
  \ee
  
\subsection{{\bf The case B:} $m_2=0.$}

The case with $m_2=L_2=0$ and again real $\Omega,\, L_1, \, g(\vec x)$ will be considered below. For such a choice, Eq.(\ref{07-3}) gives:
\be 
v_2(\vec x)=\exp(g(\vec x)/\Omega)=\frac{C}{s_2(x_2)-s_1(x_1)}, \quad v_1(\vec x)=L_1\frac{s_1(x_1)+s_2(x_2)}{s_2(x_2)-s_1(x_1)},\label{014-3}
\ee
and Eq.(\ref{06-5}) looks like:
\be 
(s_2^{\prime\prime}(x_2)-s_1^{\prime\prime}(x_1))(s_2(x_2)-s_1(x_1))-((s_1^{\prime}(\vec x_1))^2+(s_2^{\prime}(\vec x_2))^2)=CL_1(s_1(x_1)+s_2(x_2)).
\label{014-4}
\ee
The latter equation after differentiation is again, similarly to (\ref{09-3}), amenable to separation of variables but with zero in the r.h.s. It has two different solutions depending on the value of separation constant:
\be
s_1(x_1)=a_1x_1^2+b_1x_1+c_1; \quad s_2(x_2)=a_2x_2^2+b_2x_2+c_2.
 \label{014-6}
\ee
and
\be 
s_1^{\prime\prime}(x_1)=\lambda^2(s_1(x_1)+\omega_1);\quad s_2^{\prime\prime}(x_2)=-\lambda^2(s_2(x_2)+\omega_2) \label{014-5}
\ee
The latter one has three kinds of explicit solutions to insert into (\ref{014-3}):
\ba 
s_1(x_1)&=&\sigma_1\cosh(\lambda x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\lambda x_2)-\omega_2; \nonumber \\
s_1(x_1)&=&\sigma_1\sinh(\lambda x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\lambda x_2)-\omega_2; \nonumber \\
s_1(x_1)&=&\frac{1}{2}\sigma_1\exp(\lambda x_1)-\omega_1;\quad s_2(x_2)=\sigma_2\cos(\lambda x_2)-\omega_2; \nonumber 
\ea
with restrictions for constants, correspondingly:
\ba 
&&\omega_2^2-\omega_1^2+\sigma_1^2-\sigma_2^2=0; \nonumber\\
&&\omega_2^2-\omega_1^2-\sigma_1^2-\sigma_2^2=0; \nonumber \\
&&\omega_2^2-\omega_1^2-\sigma_2^2=0; \nonumber\\
&&\lambda^2(\omega_2-\omega_1)-CL_1=0.\nonumber
\ea
As for the polynomial solution (\ref{014-6}), a few restrictions have to be fulfilled simultaneously:
\be 
A\equiv2a_1+2a_2+CL_1=0. \label{016-3}
\ee
Finally, solution (\ref{014-6}) leads to two different opportunities for the components of $U_1(\vec x).$ The first one:
\ba 
v_1(\vec x)&=&-L_1\frac{[CL_1(c_2x_1^2-c_1x_2^2)+2(c_1^2-c_2^2)]}{{CL_1(c_2x_1^2+c_1x_2^2)+2(c_1-c_2)^2}};\label{017-4}\\
v_2(\vec x)&=&-\frac{2(c_1-c_2)}{{CL_1(c_2x_1^2+c_1x_2^2)+2(c_1-c_2)^2}};\label{017-5}\\
a_1&=&\frac{Cc_2L_1}{2(c_1-c_2)}; \quad a_2=-\frac{Cc_1L_1}{2(c_1-c_2)}; \, c_1\neq c_2,\nonumber
\ea
and the second, with $c_1=c_2=0$ and functions $s_1(x_1)=a_1x_1^2,\, s_2(x_2)=-(a_1+\frac{1}{2}CL_1)x_2^2,$ is:
\be 
v_1(\vec x)=-L_1\frac{a_1x_1^2-(a_1+\frac{CL_1}{2})x_2^2}{(a_1+\frac{CL_1}{2})x_2^2+a_1x_1^2};\quad 
v_2(\vec x)=-\frac{C}{(a_1+\frac{CL_1}{2})x_2^2+a_1x_1^2}.
\label{018-1}
\ee

\subsection{{\bf The case C:} $p+ic=0.$}

Let us consider again intertwining of two Dirac operators both with diagonal potential:
\be 
W_1(\vec x)=\left(
              \begin{array}{cc}
                w_{11}^{(1)}(\vec x) & 0\\ 
                0 & w_{22}^{(1)}(\vec x) \\
              \end{array}
            \right) \equiv \left(
              \begin{array}{cc}
                v_{1}(\vec x) & 0\\ 
                0 & v_{2}(\vec x) \\
              \end{array}
            \right); 
            \quad W_2(\vec x)= \left(
              \begin{array}{cc}
                k_1 & 0\\ 
                0 & k_2 \\
              \end{array}
            \right),
\label{001-0}
\ee
with constant elements $k_1, k_2.$ Here, matrix potential $W_1(\vec x)$ will be also identified with potential $V_2(\vec x)$ of Section 3, and the constant partner potential $W_2$ will provide solvability of the problem. This means that the case C is an alternative option in relation to cases B and C of the previous two Subsections. The difference with them is that we will take now $p+ic=0,$ i.e. variant III of Section 2 (it is clear that the variant IV can be considered analogously). In this case, the system (\ref{3-1}) - (\ref{4-4}) is expressed as:
\ba 
b_{11}&=&a_{22}+ibv_1-idk_2;\quad  b_{12}=\frac{n+ia}{2}v_1;\quad b_{21}=-pw_2+\frac{n-ia}{2}k_2; \nonumber \\
a_{12}&=&\frac{n+ia}{2}k_2=Const;\quad a_{21}=-pk_1+\frac{n-ia}{2}v_1. \nonumber 
\ea 
Matrix differential equation (\ref{1-9}) takes a form:
\ba 
i(n-ia)\partial v_1(\vec x)&=&v_1(\vec x)(a_{11}(\vec x)-ibk_1)-k_1(a_{22}(\vec x)-idk_2);\nonumber \\
v_2(\vec x)(a_{22}(\vec x)-idk_2)&=&k_2(a_{11}(\vec x)-ibk_1);\nonumber \\
\partial a_{22}(\vec x)&=&0;\nonumber \\
2i\bar\partial a_{11}(\vec x)&=&\frac{n-ia}{2}(v_1(\vec x)v_2(\vec x)-k_1k_2),  \nonumber
\ea 
which by convenient definition,
\be 
r_1(\vec x)\equiv a_{11}(\vec x)-ibk_1;\quad r_2(\vec x)\equiv a_{22}(\vec x)-idk_2,  \nonumber
\ee 
can be reduced to two equations $(r_2(\vec x)=r_2(\bar z))$:
\ba 
&&ik_2(n-ia)\partial (\frac{v_1(\vec x)}{r_2(\bar z)})=v_1(\vec x)v_2(\vec x)-k_1k_2; \label{003-1}\\
&&4i(\bar\partial r_1(\vec x))=(n-ia)(v_1(\vec x)v_2(\vec x)-k_1k_2). \label{003-2}
\ea 
so that both $v_1(\vec x)$ and $v_2(\vec x)$ are expressed in terms of one function $\kappa (\vec x):$
\be 
v_1(\vec x)=\frac{4r_2(\bar z)}{k_2(n-ia)}(\bar\partial \kappa (\vec x));\quad 
v_2(\vec x)=\frac{k_2(n-ia)}{r_2(\bar z)}(\partial \kappa (\vec x));\quad r_1(\vec x)=(n-ia)(\partial\kappa (\vec x)), \label{003-3}
\ee 
where function $\kappa (\vec x)$ satisfies the following second order differential equation:
\be 
i(\partial\bar\partial\kappa (\vec x))=(\partial\kappa (\vec x))(\bar\partial\kappa (\vec x))-k;\quad 4k\equiv k_1k_2. \label{003-5}
\ee
Since the r.h.s. here is real, the real part of the function $\kappa (\vec x)$ is a sum of two mutually conjugated functions:
\be 
\kappa (\vec x) \equiv \alpha (z) + \bar{\alpha} (\bar z) + i\xi (\vec x). \nonumber 
\ee 
Due to Eq.(\ref{003-3}), the reality of both diagonal elements $v_1(\vec x), \, v_2(\vec x)$ leads to reality 
of $(\partial\kappa(z,\bar z))(\bar\partial\kappa(z,\bar z))$ that means in terms of $\alpha$ and $\xi:$
\be 
\textit{Im}((\partial\kappa (\vec x))(\bar\partial\kappa (\vec x)))=\alpha^{\prime}(z)\bar\partial\xi (\vec x) + \bar\alpha^{\prime}(\bar z)\partial\xi (\vec x)=0, \nonumber 
\ee 
i.e. function $\xi (\vec x)$ is an arbitrary real function of the specific real argument:
\be 
\xi (\vec x)=\Phi (i(\alpha (z)-\bar\alpha (\bar z)))\equiv \Phi (X); \quad X(z, \bar z)\equiv i(\alpha (z)-\bar\alpha (\bar z)). \label{004-2}
\ee 
From Eq.(\ref{003-5}), we obtain nonlinear differential equation for the function $\Phi :$
\be 
\Phi^{\prime\prime}(X)-(\Phi^{\prime}(X))^2+1=\frac{k}{\alpha^{\prime}(z)\bar\alpha^{\prime}(\bar z)}. \label{004-4}
\ee
The l.h.s. in (\ref{004-4}) depends only on the variable $X$ which satisfies the following equation by definition:
\be 
(\frac{1}{\alpha^{\prime}(z)}\partial + \frac{1}{\bar\alpha^{\prime}(\bar z)}\bar\partial )X(z, \bar z)=0, \nonumber
\ee
and therefore, due to (\ref{004-4}):
\be 
(\frac{1}{\alpha^{\prime}(z)}\partial + \frac{1}{\bar\alpha^{\prime}(\bar z)}\bar\partial )\frac{1}{\alpha^{\prime}(z)\bar\alpha^{\prime}(\bar z)}=0,\nonumber 
\ee
allowing to define possible forms of the function $\alpha(z)$. Indeed, variables in the latter equation can be separated:
\be 
\frac{\alpha^{\prime\prime}(z)}{(\alpha^{\prime}(z))^2} + \frac{\bar\alpha^{\prime\prime}(\bar z)}{(\bar\alpha^{\prime}(\bar z))^2}=0 \nonumber
\ee 
informing us that exactly two options exist for the function $\alpha (z):$ 
\be 
a) \,\,\, \alpha^{\prime}(z)=\omega;\quad b)\,\,\, \alpha^{\prime}(z)=\frac{i\lambda}{z},
\label{005-00}
\ee
with $\omega $ - an arbitrary constant and $\lambda $ - an arbitrary real constant.

For the first option, $r_2(\bar z)$ must be constant, 
\ba 
v_1(\vec x)&=&4e(1+\Phi^{\prime}(X(z,\bar z)));\quad     v_2(\vec x)=\frac{|\omega|^2}{e}(1-\Phi^{\prime}(X(z,\bar z)));\label{005-3}\\ \Phi^{\prime\prime}(X)&-&(\Phi^{\prime}(X))^2+1=\frac{k}{|\omega|^2};\quad e\equiv \frac{r_2(\bar z)\bar\omega}{k_2(n-ia)}, \nonumber
\ea
and parameters must provide that the constant $e$ is real. Depending on the sign of $(\frac{k}{|\omega |^2}-1),$ one of two solutions is realized:
\be 
\Phi_1^{\prime}(X)=-\eta \cos (\eta X+\mu);\quad \Phi_2^{\prime}(X)=-\tilde\eta \ln(\cosh (\tilde\eta X+\tilde\mu));\quad (\frac{k}{|\omega |^2}-1)\equiv \eta^2>0; \,\, (\frac{k}{|\omega |^2}-1)\equiv -\tilde\eta^2<0.  \label{006-0}
\ee

For the second option, Eq.(\ref{003-3}) gives that $r_2(\bar z)=\gamma \bar z,$ constant $\frac{i\gamma }{n-ia}$ must be real, and:
\be 
v_1(\vec x) = \frac{4\lambda^2}{\beta}(1+\Phi^{\prime}(X));\quad v_2(\vec x)=\frac{\beta}{z\bar z}(1-\Phi^{\prime}(X));\quad \beta\equiv \frac{ik_2(n-ia)}{\gamma}.\label{006-1}
\ee 
Here, $\alpha (z)=i\lambda\ln (z),\,\, \bar\alpha (\bar z)=-i\lambda\ln (\bar z),$ and equation for $\Phi (X)$ takes the form:
\be 
\Phi^{\prime\prime}(X)-(\Phi^{\prime}(X))^2+1=k\lambda^2e^{\frac{X}{\lambda}}
\ee
See \cite{polyanin} about solutions of this equation in terms of special functions.

\section{Wave functions and electromagnetic fields.}

In the previous sections, we performed two consecutive transformations of Dirac operator with matrix potential using the first order intertwining operators similar to the SUSY intertwining in ordinary Quantum Mechanics. In that context, such operation is known as a second order reducible (i.e. factorizable) SUSY transformation 
\cite{AI}. Unlike that case, for the present problem with Dirac operator, the asymmetrical form of intertwining \cite{INP}, \cite{IN2022} was used in both steps. 

The resulting Dirac equation with potential which is a diagonal matrix with constant elements at the diagonal (like $U_2$ in (\ref{01-0})) is amenable to a simple analytic solution. Indeed, one of two components of $\Psi^{(2)}(\vec x)\equiv (\Psi_A^{(2)}(\vec x),\,\Psi_B^{(2)}(\vec x))^T$ can be excluded leading to the second order equation (Helmholtz equation) for another component:
\be 
(\Delta + m_1m_2)\Psi_A^{(2)}(\vec x)=0;\quad \Psi_B^{(2)}(\vec x)=\frac{2i}{m_2}\bar\partial\Psi_A^{(2)}(\vec x).
\label{A-1}
\ee
After separation of variables in (\ref{A-1}), its solution can be written as a linear combination with arbitrary complex coefficients $\sigma_{k_1k_2}:$
\be 
\Psi_A^{(2)}(\vec x)=\sum_{k_1,k_2} \sigma_{k_1k_2} \exp(k_1x_1) \exp(k_2x_2) \label{A-2}
\ee
where sum (actually, integral) is over $k_1,\,k_2$ - arbitrary complex constants such that $k_1^2+k_2^2=-m_1m_2.$ Coefficients in the sum have to be determined by boundary conditions for the wave functions.

According to the intertwining relations of the form (\ref{int}), the solutions $\Psi^{(1)}(\vec x)\equiv (\Psi_A^{(1)}(\vec x),\,\Psi_B^{(1)}(\vec x))^T$ of initial Dirac equation with potential $V_1(\vec x)$ can be constructed by sequential action of two intertwining operators $N_1\tilde N_1.$ The first operator $N_1$ intertwines two Dirac operators: initial one with potential $V_1$ and that with diagonal potential (\ref{V_2}) (see (\ref{int})). The second $\tilde N_1$ intertwines analogously Dirac operators with potential $V_2(\vec x)\equiv U_1(\vec x)\equiv W_1(\vec x)$ and operator with potential either $U_2(\vec x)$ or $W_2(\vec x),$ depending on exploring Subsection 4A,B or Subsection 4C. These intertwining operators have the general form (\ref{NM}), and the explicit expressions for the coefficients $A_k$ and $A(\vec x)$ were derived in Sections 3 and 4, correspondingly.

The two-component wave functions $\Psi^{(1)}(\vec x)$ obtained by given above procedure describe graphene and similar materials in external field: two-dimensional electrostatic plus non-homogeneous orthogonal magnetic. Analytical expressions for the strength of these fields are known from the analytical expression for initial potential $V_1(\vec x).$ The strength of electrostatic field is directed along the $(x_1, x_2)$ plane (see (\ref{A_0})):
\be 
\vec E(\vec x)=-\vec\nabla A_0(\vec x)=-\vec\nabla (f^2(\vec x)v_1(\vec x)), \label{A-3}
\ee
and the strength $B_3$ of magnetic one is (see (\ref{13-6})):
\be 
B_3(\vec x)= \triangle \ln (f(\vec x)).
\label{A-4}
\ee
Functions $f(\vec x)$ are different for different cases in Subsections 4A,B,C, and they can be calculated from the components $v_1(\vec x),\, v_2(\vec x)$ according to (\ref{13-2}). These components are given by (\ref{08-12}) and (\ref{08-22}) for {\bf 4A}; by (\ref{014-3}) for {\bf 4B}; (\ref{005-3}) and by (\ref{006-1}) for {\bf 4C}. The explicit expressions for these functions mainly in terms of trigonometric and hyperbolic functions lead to corresponding expressions for electromagnetic strengths in terms of the same elementary functions. 

For example of possible configuration of external fields we use the particular case of polynomial solution (\ref{014-6}) and (\ref{016-3}) for $c_1=c_2=0.$ The components $v_1(\vec x),\, v_2(\vec x)$ are given by (\ref{018-1}), and function $f(\vec x)$ is defined from:
\be 
f^4(\vec x)=\frac{v_2(\vec x)}{v_1(\vec x)}=\frac{2C}{L_1}[2a_1x_1^2-(2a_1+CL_1)x_2^2]^{-1}.
\nonumber 
\ee
According to (\ref{A-3}) and (\ref{A-4}), the strengths are:
\ba  
E_1(\vec x)&=&4\sqrt{2CL_1}a_1x_1\frac{2a_1x_1^2-3(2a_1+CL_1)x_2^2}{[2a_1x_1^2+(2a_1+CL_1)x_2^2]^{3}[2a_1x_1^2-(2a_1+CL_1)x_2^2]^{1/2}}; \nonumber\\ 
E_2(\vec x)&=&2\sqrt{2CL_1}(2a_1+CL_1)x_2\frac{6a_1x_1^2-(2a_1+CL_1)x_2^2}{[2a_1x_1^2+(2a_1+CL_1)x_2^2]^{3}[2a_1x_1^2-(2a_1+CL_1)x_2^2]^{1/2}}; \nonumber
\nonumber 
\ea 
and
\be 
B_3(\vec x)=2(4a_1+CL_1)(2a_1x_1^2+(2a_1+CL_1)x_2^2)(2a_1x_1^2-(2a_1+CL_1)x_2^2)^{-2}
\nonumber 
\ee 
with a possibility to choose arbitrary suitable values for all constant parameters.   

\section{Conclusions.}

The modification of well known method of SUSY Quantum Mechanics - asymmetrical intertwining relations - was used to build the massless two-dimensional Dirac equation with nontrivial matrix potential whose solutions can be found analytically. It was necessary to use factorizable second order intertwining which includes two steps: the first step allows to diagonalize the Dirac operator, and the second one - to connect the latter operator with explicitly solvable Dirac problem containing the diagonal matrix potential with constant elements. \\
{\bf Author Contributions.}

Conceptualization, writing and editing - M.V.I. and D.N.N. Both
authors have read and agreed to the published version of the
manuscript.\\
{\bf Funding.}

This research received no external funding.\\
{\bf Data Availability Statement.}

Data is contained within the article.\\
{\bf Conflicts of Interest.}

The authors declare no conflicts of interest.

\end{document}